OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN!!!! 250+ people online and no one can answer?! You owe $1,032.56 on a credit card with an 11.25% APR. the minimum due is $150. You decide to pay $300.00 instead of $150.00 at the beginning of the month. How much do you save in interest by doubling the amount?

OpenStudy (anonymous):

for just that payment or in paying off the whole amount?

OpenStudy (anonymous):

i should say for the next payment...

OpenStudy (anonymous):

i don't know. this is all the info given.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

since you pay at the beginning of the month, no interest has accrued fro the month yet. as such, you will reduce you principal by either 150 or 300. If you fgure how much interest you would pay on the 150, that is how much you will save by paying 300. (P-150)*(r/12) interest that will accumulate over the month paying 150 at the beginning (P-300)*(r/12) = (P-150-150)*(r/12) = (P-150)*(r/12) - (150)*(r/12) interest that will accumulate over the month paying 300 at the beginning (150)*(r/12) is your savings

OpenStudy (anonymous):

so \[150\cdot \frac{ .1125 }{ 12 }=$1.41\]